state whether the conclusion is coreect and why or why not

1. b=10a-60
a-6=10b
2. -2(a-3)=6
-2a-6=6
please help me I am home schooling and I am having a really hard time this algebra is tough.

Answer by Alan3354(69443)

state whether the conclusion is coreect and why or why not

1. b=10a-60
a-6=10b
-----------
Add 60
10a = b + 60
Divide by 10
a = (b/10) + 6
a - 6 = b/10, not 10b
==================
2. -2(a-3)=6
-2a-6=6
-2a + 6 = 6, not -2a - 6
---
-2 * -3 = +6